OpenStudy (anonymous):

what is 2x+2y=-6 in slope intersept form?

OpenStudy (anonymous):

I believe its 2x+2y=-6

OpenStudy (anonymous):

PLZ check the question

OpenStudy (anonymous):

yes it is equal to -6 sorry

OpenStudy (anonymous):

Slope intercept form is y=mx+c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x+2y=-6 x+y=-3 y=-x-3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yup thank you

OpenStudy (anonymous):

welcome